# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v1

@@ Abstract
Parameter-efficient fine-tuning (PEFT) techniques such as low-rank adaptation (LoRA) can effectively adapt large pre-trained foundation models to downstream tasks using only a small fraction (0.1%-10%) of the original trainable weights. An under-explored question of PEFT is in extending the pre-training phase without supervised labels; that is, can we adapt a pre-trained foundation model to a new domain via efficient self-supervised pre-training on this new domain? In this work, we introduce ExPLoRA, a highly effective technique to improve transfer learning of pre-trained vision transformers (ViTs) under domain shifts. Initializing a ViT with pre-trained weights on large, natural-image datasets such as from DinoV2 or MAE, ExPLoRA continues the unsupervised pre-training objective on a new domain. In this extended pre-training phase, ExPLoRA only unfreezes 1-2 pre-trained ViT blocks and all normalization layers, and then tunes all other layers with LoRA. Finally, we fine-tune the resulting model only with LoRA on this new domain for supervised learning. Our experiments demonstrate state-of-the-art results on satellite imagery, even outperforming fully pre-training and fine-tuning ViTs. Using the DinoV2 training objective, we demonstrate up to 7% improvement in linear probing top-1 accuracy on downstream tasks while using <10% of the number of parameters that are used in prior fully-tuned state-of-the art approaches. Our ablation studies confirm the efficacy of our approach over other baselines, including PEFT and simply unfreezing more transformer blocks.
@@ 1 Introduction
Pre-training foundation models[1] for natural language [2, 3, 4, 5] and natural images [6, 7, 8, 9, 10, 11, 12, 13] has historically been computationally intensive, often limited to organizations with substantial resources. However, recent advancements in parameter-efficient fine-tuning (PEFT) techniques including low-rank adaptation (LoRA) and others [14, 15, 16, 17, 18, 19] have sparked significant interest. These methods aim to adapt foundation models to downstream supervised-learning tasks using a small fraction (0.1%-10%) of the model’s trainable weights, based on the hypothesis that the required weight updates to the pre-trained model have a “low intrinsic rank" [14, 20, 21], or that efficient updates can be achieved by exploiting matrix structure [15, 16].
@@
In this paper, we focus on vision foundation models such as MAE or DinoV2 [7, 6], which are trained on large-scale natural-image datasets and exhibit robust generalization capabilities measured via linear probing or KNN classification for downstream tasks. Despite the large investments in developing such models for natural images, they underperform when applied to other domains with visual data (e.g., medical or remote sensing images). For example, fine-tuning a model pre-trained on natural images on satellite image classification tasks is not as effective as fine-tuning models that were pre-trained on satellite images [22, 23]. To bridge this gap, prevailing approaches invest similarly large levels of compute to pre-train foundation models on the new domains, inspired by techniques developed for natural images[22, 23, 24, 25, 26, 27].
@@
In this work, we challenge this paradigm (fig. 1), asking whether pre-training from scratch on each new domain is strictly necessary, since doing so is expensive (in compute and time) and precludes knowledge transfer from natural images. Instead, we wish to more efficiently and effectively leverage the large amount of semantic information encoded in natural-image vision foundation models to adapt them to new domains. Our proposed solution addresses these concerns using PEFT techniques for domain adaptation via self-supervised learning.
@@
Figure 1: Consider two (fairly different) image domains, $a$ and $b$. Left: the traditional paradigm of pre-training from scratch on each different domain to yield $W_{P_{a}}$ and $W_{P_{b}}$, and then fine-tuning on the target datasets $i$ to yield $\Delta_{a_{i}},\Delta_{b_{i}}$, for domains $a$ and $b$, respectively. Right: our approach, which is to initialize with pre-trained weights from domain $a$ and then learn unsupervised weights $\Delta_{P_{b}}$ for domain $b$ in a parameter-efficient manner.
@@
We introduce ExPLoRA, which generalizes vision foundation models to new domains by extending the pre-training phase with parameter-efficient techniques. We initialize a vision transformer (ViT)[28] with pre-trained weights from large, natural-image datasets (such as MAE or DinoV2). Selectively unfreezing the last, or the first and last transformer blocks, we tune remaining weights with LoRA and continue unsupervised pre-training on the new domain. Subsequently fine-tuning using linear probing or LoRA on this new domain for supervised learning outperforms prior state-of-the-art (SoTA) approaches while training less than 5%-10% of the original weights. On satellite imagery, for example, we demonstrate more than 7% improvement in linear probing top-1 accuracy, and even an improvement over prior SoTA fully pre-trained and fine-tuned techniques. We conduct an extensive study on RGB, temporal, and multi-spectral satellite images, either matching or outperforming prior methods that use full-rank pre-training from scratch. Our ablation studies confirm the superiority of our approach over alternatives. We also demonstrate generalization to different domains, such as wildlife imagery and medical images on the WILDS[29] benchmark. Our contributions include:
@@
    Introducing ExPLoRA, a novel parameter-efficient method that extends unsupervised pre-training on target domains, achieving state-of-the-art supervised-learning performance using a fraction of the original ViT weights.
@@
    Conducting a comprehensive case study on satellite imagery, showcasing improvements in linear probing top-1 accuracy and outperforming existing techniques on datasets like fMoW. We also demonstrate generalization to other domains such as wildlife and medical imagery.
@@ 2 Related Work
Visual Foundation Models
Visual foundation models(VFMs), such as DinoV2 or masked autoencoders (MAE), have demonstrated remarkable performance across downstream tasks such as classification or semantic segmentation [1, 6, 7]. However, there has also been a rise in domain-specific VFMs [22, 27, 30, 31, 32], like SatMAE, which is designed to handle temporal or multi-spectral satellite imagery[22]. With these models containing hundreds of millions of parameters, efficient adaptation to downstream tasks has become a key research focus.
@@
PEFT methods have gained widespread adoption for efficiently adapting large models to various downstream tasks, mitigating the prohibitive costs associated with full model tuning by updating only a fraction of the parameters. For example, LoRA learns low-rank weight updates to frozen weights, while other methods modify the frequency or number of trainable parameters per layer [14, 17, 18, 19, 33]. Lialin et al. [19] introduce LoRA for pre-training via the sum of multiple low-rank matrices, but require full parameter tuning as a “warm start". Some approaches use multiplicative orthogonal weight updates to frozen weights, effectively retaining pre-training knowledge [16, 15]. While our ExPLoRA method can be configured with LoRA fine-tuning for downstream tasks, it supplements existing PEFT methods rather than replacing them, particularly in the case of unsupervised domain adaptation.
@@
The central problem in domain adaptation is managing the distribution shift with respect to training and testing data. Domain adaptation approaches have explored this issue from several perspectives[34]. Discrepancy-based methods minimize the difference between feature distributions of the source and target domains using discrepancy metrics[35, 36, 37] for domain loss. Adversarial methods aim to amplify domain confusion while simultaneously being rigorously trained to recognize and distinguish between different domains [38, 39, 40]. Group DRO aims to minimize the loss in the worst-case domain within the context of domain adaptation, specifically addressing subpopulation shift, where data distributions differ but may have some overlap [41, 42].
@@ 3 Background
The masked-autoencoder (MAE) [7] is an effective self-supervised learning technique for ViTs. It utilizes an asymmetrical encoder-decoder architecture on images ${\mathbf{x}}\in{\mathbb{R}}^{C\times H\times W}$ (height, width, channels $H,W,C$), where patches are masked before being processed by the ViT encoder ${\mathcal{L}}$. The visible embedded patches are then reconstructed using a smaller ViT decoder ${\mathcal{L}}_{D}$, with the training objective being a mean-squared error loss on the visible pixels. MAE has been widely adopted in various domains including satellite imagery [22], video [43], and other modalities[44, 45]. However, pre-training ViTs as MAEs typically requires subsequent full fine-tuning for downstream tasks, making them computationally intensive.
@@
DinoV2 [6] is a robust self-supervised learning technique for ViTs. Unlike MAE, DinoV2 features have demonstrated strong zero-shot performance, enabling adaptation to downstream tasks even with a frozen ViT backbone. During pre-training, DinoV2 maintains two copies of a ViT encoder: the student (trainable) and the teacher, which is updated using an exponential-moving average of the student’s parameters. The training objective incorporates a global, image-level loss from Dino [11] and a patch-based loss from iBOT [8]. Various techniques including KoLeo regularization [46] and Sinkhorn-Knopp centering [10] are utilized, as described in [6].
@@
Low-rank adaptation (LoRA) [14] assumes that the weight update to change a set of unsupervised pre-trained weights to supervised fine-tuned weights lives in a low-rank subspace,
$W\approx W_{0}+\Delta_{W}=W+BA$ (1)
@@ 4 Problem Setup
Consider a set of image domains ${\mathcal{D}}=\{1,2,\dots\}$ and an associated data distribution for each domain $p_{d}({\mathbf{x}})$, where $d\in{\mathcal{D}}$ and images ${\mathbf{x}}\in{\mathbb{R}}^{C_{d}\times H_{d}\times W_{d}}$. Here $C_{d}$, $H_{d}$ and $W_{d}$ represent the channel, height and width dimensions of images of each domain, respectively. Let $D_{P},D_{F}\subset{\mathcal{D}}$ be a set of domains representing the pre-training and fine-tuning data distributions $p_{D_{P}}({\mathbf{x}})$ (eg: internet-scale natural image data) and $p_{D_{F}}({\mathbf{x}})$ (eg: satellite data), respectively. Next, the fine-tuning joint distributions for each domain $d_{F}\in D_{F}$ are $p_{d_{F}}({\mathbf{x}},{\mathbf{y}})$, where ${\mathbf{y}}$ is the supervised-learning label.
We then assume access to the following: (i) pre-trained weights $W_{D_{P}}$, indexed by the collection of pre-training domains $D_{P}$, which have already been obtained via unsupervised learning (ii) samples from $p_{D_{F}}({\mathbf{x}})$ representing unlabeled images from a new, different domain $D_{F}$ (iii) a collection of target datasets $d_{F}\in D_{F}$ from the new domain $D_{F}$, sampled from distributions $p_{d_{F}}({\mathbf{x}},{\mathbf{y}})$. Thus, we would like to learn optimal weights $W_{d_{F}}$ in a parameter-efficient manner for each supervised-learning dataset while leveraging knowledge stored in $W_{D_{P}}$.
LoRA solves this in the following way:
$W_{d_{F}}\approx W_{D_{P}}+\Delta_{d_{F}}=W_{D_{P}}+B_{d_{F}}A_{d_{F}}$ (2)
The LoRA hypothesis is that the update $\Delta_{d_{F}}$ resides in a low-rank subspace when adapting pre-trained weights $W_{D_{P}}$ to fine-tuned weights $W_{d_{F}}$. This hypothesis empirically seems to work when pre-training and fine-tuning distributions are similar, or where $d_{F}\in D_{P}$. However, we empirically find that when adapting to a significantly different domain, where $D_{F}$ has little to no overlap with $D_{P}$ (eg: natural images to multi-spectral satellite data), this hypothesis fails to hold (see section 6.1.3).
@@
To address this challenge, the traditional solution (fig. 1) has been to begin pre-training from scratch on the new domains of interest $D_{F}$, and then use PEFT for each dataset $d_{F}$, representing the following:
$W_{d_{F}}\approx W_{D_{F}}+\Delta_{d_{F}}$ (3)
where $W_{D_{F}}$ now represent unsupervised pre-trained weights learned from the data distribution $p_{D_{F}}({\mathbf{x}})$. However, this is wasteful: learning $W_{D_{F}}$ from scratch can require prohibitively large amounts of additional compute to complete full-rank pre-training for each new domain of interest.
Our goal is thus the following: can we still leverage the rich semantic information encoded in $W_{D_{P}}$ and learn $W_{D_{F}}$ for a new domain in a parameter-efficient manner? Concretely, we would like to factorize $W_{d_{F}}$ as follows:
$W_{d_{F}}\approx W_{D_{P}}+\Delta_{D_{F}}+\Delta_{d_{F}}$ (4)
where $\Delta_{D_{F}}\in\mathbb{R}^{k_{2}\times k_{1}}$ is an additional update matrix learned for the collection of fine-tuning domains $D_{F}$. Our key requirements for $\Delta_{D_{F}}$ are: (i) $\Delta_{D_{F}}$must be learned via unsupervised pre-training on $p_{D_{F}}({\mathbf{x}})$ (ii) $\Delta_{D_{F}}$must only require learning a fraction of the $k_{1}k_{2}$ parameters that form $W_{D_{P}}$
Note that successfully learning $\Delta_{D_{F}}$, would obviate the vast computing resources that are otherwise necessary to fully train foundation models for the new domain. Importantly, we emphasize learning $\Delta_{D_{F}}$ in an unsupervised manner so that the resulting model $W_{D_{F}}^{\prime}=W_{D_{P}}+\Delta_{D_{F}}\approx W_{D_{F}}$ retains the benefits of pre-trained foundation models, such as feature extraction, effective linear-probing, and generalization to further downstream tasks.

### v2

@@ Abstract
Parameter-efficient fine-tuning (PEFT) techniques such as low-rank adaptation (LoRA) can effectively adapt large pre-trained foundation models to downstream tasks using only a small fraction (0.1%-10%) of the original trainable weights. An under-explored question of PEFT is in extending the pre-training phase without supervised labels; that is, can we adapt a pre-trained foundation model to a new domain via efficient self-supervised pre-training on this new domain? In this work, we introduce ExPLoRA, a highly effective technique to improve transfer learning of pre-trained vision transformers (ViTs) under domain shifts. Initializing a ViT with pre-trained weights on large, natural-image datasets such as from DinoV2 or MAE, ExPLoRA continues the unsupervised pre-training objective on a new domain, unfreezing 1-2 pre-trained ViT blocks and tuning all other layers with LoRA. We then fine-tune the resulting model only with LoRA on this new domain for supervised learning. Our experiments demonstrate state-of-the-art results on satellite imagery, even outperforming fully pre-training and fine-tuning ViTs. Using the DinoV2 training objective, we demonstrate up to 7.5% improvement in linear probing top-1 accuracy on downstream tasks while using <10% of the number of parameters that are used in prior fully-tuned state-of-the art approaches. Our ablation studies confirm the efficacy of our approach over other baselines, including PEFT and unfreezing more ViT blocks. Code is available on the project website: https://samar-khanna.github.io/ExPLoRA/
@@ 1 Introduction
Pre-training foundation models (Bommasani et al., 2021) for natural language (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023; Jiang et al., 2024) and natural images (Oquab et al., 2023; He et al., 2022; Zhou et al., 2021; Rombach et al., 2022) has historically been computationally intensive, often limited to organizations with substantial resources. However, recent advancements in parameter-efficient fine-tuning (PEFT) techniques including low-rank adaptation (LoRA) and others (Hu et al., 2021; Zhang et al., 2023b; Chavan et al., 2023; Qiu et al., 2023; Liu et al., 2023; Jia et al., 2022) have sparked significant interest. These methods aim to adapt foundation models to downstream supervised-learning tasks using a small fraction (0.1%-10%) of the model’s trainable weights, with many based on the hypothesis that the required weight updates to the pre-trained model have a “low intrinsic rank" (Hu et al., 2021; Li et al., 2018a; Aghajanyan et al., 2020).
@@
In this paper, we focus on visual foundation models (VFMs) such as DinoV2 or MAE (Oquab et al., 2023; He et al., 2022), which were trained on large-scale natural-image datasets. Despite the large investments in developing such models for natural images, they underperform when applied to other domains with visual data (e.g. medical or satellite images). For example, fine-tuning a model pre-trained on natural images on satellite image classification tasks is not as effective as fine-tuning one that was pre-trained on satellite images (Cong et al., 2022; Ayush et al., 2021). To bridge this gap, prevailing approaches invest similarly large levels of compute to pre-train VFMs on new domains, inspired by techniques developed for natural images (Cong et al., 2022; Reed et al., 2023; Tang et al., 2024; Khanna et al., 2024; Zhou et al., 2023; Moutakanni et al., 2024; Man et al., 2023).
@@
In this work, we challenge this paradigm (fig. 1), asking whether pre-training from scratch on each new domain is strictly necessary, since doing so is expensive (in compute and time) and precludes knowledge transfer from natural images. Instead, we wish to more efficiently leverage the rich semantic information encoded in natural-image vision foundation models to adapt them to new domains. Our proposed solution addresses these concerns using PEFT techniques for domain adaptation via self-supervised learning.
@@
Figure 1: Consider two different image domains, $D_{S}$ and $D_{T}$. Left: the traditional paradigm of pre-training from scratch on each domain to yield $W_{D_{S}}$ and $W_{D_{T}}$, and then fine-tuning on target datasets $i$ to yield $\Delta_{s_{i}},\Delta_{t_{i}}$, for domains $D_{S}$ and $D_{T}$, respectively. Right: our approach, which is to initialize with pre-trained weights from domain $D_{S}$ and then learn unsupervised weights $\Delta_{D_{T}}$ for domain $D_{T}$ in a parameter-efficient manner.
@@
We introduce ExPLoRA, which generalizes vision foundation models to new domains by extending the pre-training phase with parameter-efficient techniques. We initialize a vision transformer (ViT) (Dosovitskiy et al., 2021) with pre-trained weights from natural-image datasets such as MAE or DinoV2. Selectively unfreezing 1-2 transformer blocks, we tune remaining weights with LoRA and continue unsupervised pre-training on the new domain. Subsequently fine-tuning with linear probing or LoRA on this new domain for supervised learning outperforms prior state-of-the-art (SoTA) approaches while training under 6-10% of the original weights. On satellite imagery, for example, we demonstrate an 8% improvement in linear probing top-1 accuracy, and even an improvement over prior SoTA fully pre-trained and fine-tuned techniques. We conduct an extensive study on RGB, temporal, and multi-spectral satellite images, either matching or outperforming prior methods that fully pre-train from scratch. ExPLoRA also generalizes to different domains such as wildlife, medical, and agricultural imagery on the WILDS (Koh et al., 2021) benchmark. Our contributions include:
@@
    Introducing ExPLoRA, a novel parameter-efficient method that extends unsupervised pre-training on target domains, achieving state-of-the-art supervised-learning performance using a fraction of the original ViT weights (section 5).
@@
    Conducting a comprehensive case study on satellite imagery, showcasing improvements in linear probing top-1 accuracy and outperforming existing techniques on datasets like fMoW. We also demonstrate generalization to multiple other domains within WILDS (section 6).
  3. 3.
    Demonstrating ExPLoRA’s efficacy via ablation studies and by analyzing the differences in local (eg: positional) and global (eg: class) information encoded in the patch representations output by each ViT block (section 6.3).
@@ 2 Related Work
VFMs
VFMs such as DinoV2 or masked autoencoders (MAE) that pre-train with self-supervised learning (SSL) have demonstrated remarkable performance across downstream tasks such as classification or semantic segmentation Oquab et al. (2023); He et al. (2022); Grill et al. (2020); Chen et al. (2020). However, there has also been a rise in domain-specific VFMs (Cong et al., 2022; Reed et al., 2023; Moutakanni et al., 2024; Ma et al., 2024; Zhang et al., 2023a). For instance, SatMAE handles temporal or multi-spectral satellite image inputs. Since these models contain hundreds of millions of parameters, efficient adaptation to downstream tasks has become a key research focus.
@@
PEFT methods have gained widespread adoption for efficiently adapting large models to various downstream tasks, mitigating the prohibitive costs of full model tuning by updating only a fraction of the parameters. For example, LoRA learns low-rank weight updates to frozen weights, while other methods modify the frequency or number of trainable parameters per layer (Hu et al., 2021; Zhang et al., 2023b; Chavan et al., 2023; Pu et al., 2023). Lialin et al. (2023) introduce LoRA for pre-training via adding multiple low-rank matrices, but require full parameter tuning as a “warm start”. Others use multiplicative orthogonal weight updates, effectively retaining pre-training knowledge in frozen weights (Qiu et al., 2023; Liu et al., 2023). Another family of PEFT techniques include visual prompt tuning (VPT), which concatenate learnable prompt tokens to the sequence of image patch tokens, trading improved fine-tuning performance with increased inference-time latency and memory usage (Jia et al., 2022; Yoo et al., 2023; Pei et al., 2024; Han et al., 2023). Our ExPLoRA method aims to supplement existing PEFT methods rather than replace them, and thus can be configured with any existing or future PEFT method intended for fine-tuning ViT backbones.
@@
Domain adaptation approaches have attempted to bridge the train-test data distribution shift from several perspectives (Singhal et al., 2023). Discrepancy-based methods minimize the difference between feature distributions of the source and target domains using discrepancy metrics (Gretton et al., 2012; Kang et al., 2019; Sun et al., 2016) for domain loss. Adversarial methods are trained to amplify domain confusion while simultaneously distinguishing between different domains (Ganin et al., 2016; Li et al., 2018b; Tzeng et al., 2017). Group DRO minimizes the loss in the worst-case domain, specifically addressing subpopulation shift, where data distributions differ but may have some overlap (Hu et al., 2018; Oren et al., 2019).
@@ 3 Background
The masked-autoencoder (MAE) (He et al., 2022) is an effective SSL technique for ViTs that uses an asymmetric encoder-decoder architecture on images ${\mathbf{x}}\in{\mathbb{R}}^{C\times H\times W}$, where patches are masked before being processed by the ViT encoder ${\mathcal{L}}$. The masked patches are then reconstructed by a smaller decoder ${\mathcal{L}}_{D}$, with both trained jointly using mean-squared error on the reconstructed visible pixels. While effective across domains (Cong et al., 2022; Bachmann et al., 2022), MAEs typically require full fine-tuning for downstream tasks, which makes them computationally expensive.
@@
DinoV2 (Oquab et al., 2023) is a robust SSL method for ViTs. Unlike MAE, DinoV2 features have demonstrated strong zero-shot performance, enabling adaptation to downstream tasks even with a frozen ViT backbone. During pre-training, DinoV2 maintains two copies of a ViT encoder: the student (trainable) and the teacher, which is updated using an exponential-moving average of the student’s parameters. The training objective incorporates a global, image-level loss from Dino (Caron et al., 2021), a patch-based loss from iBOT (Zhou et al., 2021), and regularizers including KoLeo (Delattre & Fournier, 2017) and Sinkhorn-Knopp centering (Caron et al., 2020).
@@
Low-rank adaptation (LoRA) (Hu et al., 2021) assumes that the weight update to change a set of unsupervised pre-trained weights to supervised fine-tuned weights lives in a low-rank subspace,
$W\approx W_{0}+\Delta_{W}=W_{0}+BA$ (1)
@@ 4 Problem Setup
Consider a set of image domains ${\mathcal{D}}=\{1,2,\dots\}$, where each domain $d\in{\mathcal{D}}$ is associated with a data distribution $p_{d}({\mathbf{x}})$, and images ${\mathbf{x}}\in{\mathbb{R}}^{C_{d}\times H_{d}\times W_{d}}$ have domain-specific channel, height, and width. Let $D_{S}\subset{\mathcal{D}}$ represent a set of source domains (e.g., internet-scale natural image data) and $D_{T}\subset{\mathcal{D}}$ represent target domains (e.g., satellite imagery). The data from the source domains follow a distribution $p_{D_{S}}({\mathbf{x}})$, and the target domain data come from $p_{D_{T}}({\mathbf{x}})$. For some target domains $d_{T}\in D_{T}$, the joint distributions $p_{d_{T}}({\mathbf{x}},{\mathbf{y}})$ describe images ${\mathbf{x}}$ with associated supervised labels ${\mathbf{y}}$ used for downstream tasks. We then assume access to the following:
  1. (i)
    $W_{D_{S}}$, pre-trained weights obtained via unsupervised pre-training on images from $p_{D_{S}}({\mathbf{x}})$
  2. (ii)
    ${\mathcal{X}}_{D_{T}}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\sim p_{D_{T}}({\mathbf{x}})$, an unlabeled dataset of $N$ images from new domains $D_{T}$
  3. (iii)
    ${\mathcal{Y}}_{d_{T}}=\{{\mathbf{x}}_{j},{\mathbf{y}}_{j}\}_{j=1}^{M_{d_{T}}}\sim p_{d_{T}}({\mathbf{x}},{\mathbf{y}})$ a labeled dataset of $M_{d_{T}}$ images from domain $d_{T}\in D_{T}$
Our objective is to learn optimal weights $W_{d_{T}}$ for each supervised-learning dataset ${\mathcal{Y}}_{d_{T}}$ in a parameter-efficient manner while leveraging the knowledge stored in $W_{D_{S}}$.
@@
Traditionally, the approach (fig. 1) has been to begin pre-training from scratch on the new domains of interest in ${\mathcal{X}}_{D_{T}}$, and then fine-tune for each dataset ${\mathcal{Y}}_{d_{T}}$, representing the following:
$W_{d_{T}}\approx W_{D_{T}}+\Delta_{d_{T}}$ (2)
where $W_{D_{T}}$ represents the weights learned from unsupervised pre-training on ${\mathcal{X}}_{D_{T}}$, and $\Delta_{d_{T}}$ are the weights learned from supervised fine-tuning on ${\mathcal{Y}}_{d_{T}}$. However, this method is computationally expensive: fully pre-training $W_{D_{T}}$ from scratch for every new domain requires prohibitively large amounts of additional compute.
On the other hand, LoRA addresses this inefficiency in the following way:
$W_{d_{T}}\approx W_{D_{S}}+\Delta_{d_{T}}=W_{D_{S}}+B_{d_{T}}A_{d_{T}}$ (3)
The LoRA hypothesis is that the update $\Delta_{d_{T}}$ resides in a low-rank subspace when adapting pre-trained weights $W_{D_{S}}$ to fine-tuned weights $W_{d_{T}}$. This hypothesis holds well when pre-training and fine-tuning distributions are similar, or where $d_{T}\in D_{S}$. However, when there is significant domain shift, such as between natural images and multi-spectral satellite data, the low-rank assumption often breaks down (see section 6.1.3).
Our goal is to learn $W_{D_{T}}$ in a parameter-efficient manner to bridge the large domain shift to $D_{T}$ while leveraging the knowledge encoded in $W_{D_{S}}$. We propose the following factorization of $W_{d_{T}}$:
$W_{d_{T}}\approx W_{D_{S}}+\Delta_{D_{T}}+\Delta_{d_{T}}$ (4)
where $\Delta_{D_{T}}\in\mathbb{R}^{k_{2}\times k_{1}}$ is an additional update matrix learned from unsupervised pre-training on ${\mathcal{X}}_{D_{T}}$. Crucially, $\Delta_{D_{T}}$ requires only a fraction of the $k_{1}k_{2}$ parameters of $W_{D_{S}}$, making it significantly more efficient than full-rank pre-training. The resulting model, $W_{D_{T}}^{\ast}=W_{D_{S}}+\Delta_{D_{T}}\approx W_{D_{T}}$, retains the benefits of unsupervised pre-trained VFMs, including strong feature extraction, effective linear probing, KNN classification, and generalization to downstream tasks.
Algorithm 1 ExPLoRA
[…]
2:Initialize a frozen ViT with $W_{D_{S}}$ from source domains $D_{S}$ (e.g., DinoV2 or MAE weights).
[…]
4:Apply LoRA (with rank $r$) on $Q$ and $V$ weights in attention layers of frozen blocks in ${\mathcal{L}}\setminus{\mathcal{U}}$ and unfreeze normalization layers in these blocks.
5:Train all unfrozen parameters $\Delta_{D_{T}}$ on the unlabeled dataset ${\mathcal{X}}_{D_{T}}$ using the same unsupervised objective as what was used for $W_{D_{S}}$ (e.g., DinoV2 or MAE).
6:Output: A new pre-trained model $W_{D_{T}}^{\ast}=W_{D_{S}}+\Delta_{D_{T}}$ for target domains $D_{T}$.